# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ 3.2 Adaptively estimating a good flippancy bound $w$
In this section, we leverage the privacy and accuracy guarantees of Algorithm 1 to construct a new mechanism (Algorithm 3) for estimating $\mathsf{CountDistinct}$ that achieves the privacy and accuracy guarantees of Theorem 3.1, when the maximum flippancy is not known upfront. Algorithm 3 instantiates $\log T+1$ different copies $\mathcal{B}_{0},\dots\mathcal{B}_{\log T}$ of Algorithm 1 with flippancy bounds $2^{0},\dots,2^{\log T}$, respectively (the maximum flippancy of any stream is necessarily bounded by $T$). To obtain an accurate estimate of the distinct elements count, at each time $t\in[T]$, we privately select $i\in[0,\log T]$ such that the output of $\mathcal{B}_{i}$ satisfies the desired accuracy guarantee for the stream entries $x[1:t]$ received so far. Selecting such $i$ amounts to selecting a good bound on the maximum flippancy of the stream $x[1:t]$. Next, we describe how to obtain this bound using the sparse vector technique (Algorithm 2).
Event-level privacy.
For event-neighboring streams, the maximum flippancy has sensitivity 1: changing one entry in the stream can change the maximum flippancy by at most 1. Hence, we can use the sparse vector technique (Algorithm 2) to monitor the maximum flippancy of the stream — in particular we monitor when the maximum flippancy of the stream has doubled.
Item-level privacy.
For this case, the maximum flippancy is no longer a sensitivity-one function; changing all entries in the stream pertaining to one item can change the maximum flippancy drastically. However, the number of items with flippancy greater than any particular threshold is still a function of sensitivity one. Furthermore, since Algorithm 1 when run with flippancy bound $w$ already has error $O(\sqrt{w/\rho})$, its accuracy guarantee remains the same even if it simply ignores that many elements with flippancy greater than $w$. Thus, Algorithm 3 uses the sparse vector technique to maintain an upper bound on the flippancy of $x[1:t]$ such that not too many elements in $x[1:t]$ violate that bound. This bound, in combination with the error guarantee of Algorithm 1, suffices to provide the desired low error guarantee. Since the sparse vector technique (Algorithm 2) remains differentially private even when its queries are chosen adaptively, the privacy guarantees of Algorithm 3 follow from the privacy of Algorithms 1 and 2.
@@
8:         Output $\mathsf{Above}$
@@
11:         Output $\mathsf{Below}$
@@
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, stream $x\in\mathcal{U}_{\pm}^{T}$.
@@
10:     for all $i\in[\log{T}]$ do
11:         Send $x[t]$ to mechanism $\mathcal{B}_{i}$ and get output $r_{i,t}$
@@
16:     Output $s_{j,t}$ for $j=\log(w_{\max}[t])$triangleright Note that $j\in[0,\log T]$
@@
Algorithm 2 is $\rho$-zCDP. Let $k$ be the index of the last “$\mathsf{Above}$” query answered by Algorithm 2 (before cutoff $c$ has been crossed). With probability at least $1-\beta$, the vector of answers to the queries $q_{1},\dots,q_{k}$ is $\gamma$-accurate for $\gamma=\frac{8c(\ln k+\ln(2c/\beta))}{\sqrt{2\rho}}$.
@@
We start by showing that Algorithm 3 is $\rho$-item-level-zCDP. Algorithm 3 accesses the stream $x$ via the algorithms $\mathcal{B}_{i}$, $i\in[0,\log T]$ (Algorithm 1) and Algorithm 2. We showed in Theorem 3.2 that Algorithm 1 with privacy parameter $\rho^{\prime}$ is $\rho^{\prime}$-item-level-zCDP. Since we use $(\log T+1)$ instantiations of Algorithm 1, each with privacy parameter $\frac{\rho}{2(\log T+1)}$, by composition, the aggregate of the calls to Algorithm 1 is $(\frac{\rho}{2})$-item-level-zCDP. We now show that the aggregate of the calls to Algorithm 2 is $(\frac{\rho}{2})$-item-level-zCDP. Note that the queries $q_{t}$ for $t\in[T]$ considered in Step 13 of Algorithm 3 have sensitivity $1$ for item-neighboring streams (the number of items with flippancy above a certain threshold can change by at most 1 for item-neighboring streams). By Theorem 3.7, the calls to Algorithm 2 are $(\frac{\rho}{2})$-item-level-zCDP. Another invocation of the composition lemma gives that Algorithm 3 is $\rho$-item-level-zCDP.
We now analyze the accuracy of Algorithm 3. Set $\beta_{\mathsf{SVT}}=0.005$, $k=T$, $c=\log T$, and $\gamma_{\mathsf{SVT}}=\frac{8\log T(\log T+\log(400\log T))}{\sqrt{2\rho}}$. Let $E$ be the event that the vector of answers output by the sparse vector algorithm (Algorithm 2) until the cutoff point $\log T$ is $\gamma_{\mathsf{SVT}}$-accurate. By Theorem 3.7, the probability of $E$ is at least $0.995$. We condition on $E$ for most of the following proof.
@@
Let $w_{t}$ be the (true) maximum flippancy of the sub-stream $x[1:t]$, consisting of the first $t$ entries of the input stream $x\in\mathcal{U}_{\pm}^{T}$ to Algorithm 3. Then, for all $t\in[T]$, when the algorithm reaches Step 16, we have
@@
Let $z$ be the value of $w_{\max}[t]$ when Algorithm 3 reaches Step 16. If $z=1$ then $z=w_{\max}[t]\leq 2\gamma^{2}_{\mathsf{SVT}}$ since $T>1,\rho<1$. So, instead assume that $z\geq 2$. Let $t^{*}\leq t$ be the time step where $w_{\max}[t^{*}]$ is doubled from $\frac{z}{2}$ to $z$ during an execution of Step 15 of the $\mathbf{while}$ loop. This only happens if Algorithm 2 outputs ”$\mathsf{Above}$” for the following query:
@@
Suppose there is some $t\in[T]$ during which $\mathsf{count}\geq c$. Consider the last time step $t^{*}\in[T]$ when Step 9 of Algorithm 2 is run (for this time step, $\mathsf{count}=c-1$). At this time step, $w_{max}[t^{*}]$ doubles from $\frac{T}{2}$ to $T$, after which it never changes again. By case (1), we have that $w_{max}[t^{*}]=T\leq\max(2w_{t^{*}},2\gamma^{2}_{\mathsf{SVT}})$. Since for all $t\geq t^{*}$ it holds $w_{t^{*}}\leq w_{t}$ and $w_{max}[t]=w_{max}[t^{*}]$, then $w_{max}[t]\leq\max(2w_{t},2\rho\gamma^{2}_{\mathsf{SVT}})$ for all $t\geq t^{*}$.
This concludes the proof of Claim 3.9. ∎
Now we substitute the upper bound on $w_{max}[t]$ from Claim 3.9 into Equation (4) and use the fact that $w_{t}\leq w$. We get that, for sufficiently large $T$, conditioned on event $E$, with probability at least $0.995$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is
$\operatorname{O}\left(\frac{\log^{2}T+\sqrt{\max(w,2\rho\gamma^{2}_{\mathsf{SVT}})\log^{2}T}}{\sqrt{\rho}}\right)$
@@
Finally, by a union bound over the above failure probability and that of $E$: For sufficiently large $T$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is bounded by Equation (5) with probability at least $0.99$. ∎
@@
Let $f\colon\mathcal{U}_{\pm}^{*}\to\mathbb{R}$ be a function of $\ell_{2}$-sensitivity at most $1$. Define $F\colon\mathcal{U}_{\pm}^{T}\to\mathbb{R}^{T}$ so that $F(x)=[f(x[1:1]),\dots,f(x[1:T])]$. Fix $\rho\in(0,1]$ and sufficiently large $T\in N$. Then, there exists a $\rho$-item-level-zCDP mechanism for estimating $F$ in the continual release model that is $\alpha$-accurate where $\alpha=\operatorname{O}\left(\min\left\{\sqrt[3]{\frac{T\log T}{\rho}},T\right\}\right).$
Note that $\mathsf{CountDistinct}(x)[t]$ has $\ell_{2}$-sensitivity one for item-neighboring streams for all $t\in[T]$. Let $\mathcal{M}^{\prime}$ be the mechanism from Theorem 3.10. Then $\mathcal{M}^{\prime}$ can be used for estimating $\mathsf{CountDistinct}$ under continual release for turnstile streams with the error guarantee stated in Theorem 3.10. When the maximum flippancy of the stream is larger than roughly $\rho^{1/3}T^{2/3}$, the mechanism $\mathcal{M}^{\prime}$ achieves better error than that of Theorem 3.1 (and it achieves worse error when the maximum flippancy of the stream is smaller than this threshold). A simple modification of Algorithm 3 can get the best of both worlds–instead of having base mechanisms $B_{0},\dots,B_{\log T}$ that each run Algorithm 1 with different flippancy parameters as input, we only have $k+2=\mathsf{min}(\rho^{1/3}T^{2/3},T)$ base mechanisms $B_{0},\dots,B_{k+1}$. Out of these, $B_{0},\dots,B_{k}$ run Algorithm 1, whereas $B_{k+1}$ runs $\mathcal{M}^{\prime}$. A similar proof to that of Theorem 3.1, which uses the error guarantee of the recompute- mechanism $\mathcal{M}^{\prime}$ for base mechanism $B_{k+1}$ instead of the guarantee of Algorithm 1, gives an error upper bound of $O\left(\mathsf{min}\left(\sqrt{\frac{w}{\rho}}\operatorname{polylog}T,\sqrt[3]{\frac{T\log T}{\rho}},T\right)\right)$.
@@ 4 Event-level privacy lower bound
In this section, we prove Theorem 1.5 that provides strong lower bounds on the accuracy parameter $\alpha$ of any accurate, event-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
@@ 4.1 Reduction from $\mathsf{InnerProducts}$
We obtain our lower bound by showing that any mechanism for $\mathsf{CountDistinct}$ for turnstile streams can be used to obtain algorithms with similar accuracy guarantees for $\mathsf{InnerProducts}$, the problem of estimating answers to inner product queries. The reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$ combines two ideas: one is the sequential embedding technique introduced by Jain et al. [43] to prove lower bounds in the continual release model and the other is a connection between the inner product of two vectors and the number of distinct elements in their union. The latter idea was used by Mir et al. [49] to give lower bounds for pan-private algorithms counting the number of distinct elements. The reduction is presented in Algorithm 4. With this reduction, we can then use previously established lower bounds on accuracy for $\mathsf{InnerProducts}$ ([21, 24, 49, 18]) to obtain our lower bound on $\mathsf{CountDistinct}$. We start by proving Lemma 4.2 (the reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$). In Section 4.2, we use Lemma 4.2 to complete the proof of Theorem 1.5
@@
2:Output: Estimates of inner product queries $b=(b[1],\dots,b[k])\in\mathbb{R}^{k}$
3:Define the universe $\mathcal{U}=[n]$.
4:Let $x^{(1)}$ be a stream of length $n$
@@
6:     If $y[i]=1$ set $x^{(1)}[i]=+i$, otherwise set $x^{(1)}[i]=\perp$
7:Intialize streams $z^{(1)}=\perp^{2n},\dots,z^{(k)}=\perp^{2n}$
@@
9:     Set $z^{(j)}[i]=+i$triangleright phase one
10:     Set $z^{(j)}[n+i]=-i$triangleright phase two
11:Run $\mathcal{M}$ on the stream $x\leftarrow x^{(1)}\circ z^{(1)}\circ z^{(2)}\circ\dots\circ z^{(k)}$ and record the answers as vector $r$ of length $(2k+1)n$
12:for all $j\in[k]$ do
13:     Compute $\|q^{(j)}\|_{0}$ and let $b[j]=\|q^{(j)}\|_{0}+r[n]-r[2jn]$
14:Output the estimates $(b[1],\dots,b[k])$
@@
Let $\mathcal{A}$ be Algorithm 4. For all $\varepsilon>0,\delta\geq 0,\alpha\in\mathbb{R}^{+}$ and $n,T,k\in\mathbb{N}$, where $T\geq(2k+1)n$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-event-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $2k$, then batch algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$.
The proof of Lemma 4.2 crucially uses the following connection between the inner product of two vectors and the number of distinct elements in their union.
Definition 4.3 (Stream Indicator).
Remark 4.4.
For any two insertion-only streams $x$ and $x^{\prime}$, the following holds:
With this remark, we can now prove Lemma 4.2.
@@
Event-level differential privacy follows from the fact that $\mathcal{M}$ is event-level-DP and changing a record of the dataset $y$ corresponds to changing a single entry of the stream $x$, and more specifically, an entry of the stream $x^{(1)}$ constructed in Step 6 of Algorithm 4.
So we are left to prove the accuracy of $\mathcal{A}$. Fix queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$. Firstly, observe that for all $j\in[k]$ , the stream $z^{(j)}$ is constructed so that $q^{(j)}$ is the indicator vector for $z^{(j)}[1:n]$, namely the first half of $z^{(j)}$. Similarly, $x^{(1)}$ is constructed such that $y$ is its stream indicator vector.
Next, since at time $2jn$, all of the stream entries pertaining to earlier queries $q^{(1)},\dots,q^{(j-1)}$ have been deleted and those pertaining to $q^{(j)}$ have been added, we have that $\|h_{x[1:2jn]}\|_{0}=\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}$, for $j\in[k]$.
The streams $x^{(1)}$ and $z^{(j)}[1:n]$ for $j\in[k]$, are all insertion-only streams. Hence, by Remark 4.4,
$\langle h_{x^{(1)}},h_{z^{(j)}[1:n]}\rangle=\|h_{x^{(1)}}\|_{0}+\|h_{z^{(j)}[1:n]}\|_{0}-\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}.$
As observed earlier, $h_{x^{(1)}}=y$, $h_{z^{(j)}[1:n]}=q^{(j)}$, and $\|h_{x[1:2jn]}\|_{0}=\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}$, which gives that
$\langle y,q^{(j)}\rangle$ $=\|h_{x^{(1)}}\|_{0}+\|q^{(j)}\|_{0}-\|h_{x[1:2jn]}\|_{0}$ (6)
@@
Since the mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ is $\alpha$-accurate on the constructed stream then, with probability at least $0.99$, the answers of $\mathcal{M}$ are within additive error $\alpha$ of the distinct counts of the corresponding stream prefixes. Condition on this event for the rest of this proof. Then, $|r[n]-\|h_{x^{(1)}}\|_{0}|\leq\alpha$. Similarly, $|r[2jn]-\|h_{x[1:2jn]}\|_{0}|\leq\alpha$ for all $j\in[k]$. Additionally, $\|q^{(j)}\|_{0}$ is computed exactly by $\mathcal{A}$. Hence, by the triangle inequality, Equation 6, and the setting of $b[j]$ in Step 13, we have that $|b[j]-\langle y,q^{(j)}\rangle|\leq 2\alpha$. This argument applies for all $j\in[k]$. Hence, with probability at least $0.99$, all of the estimates $b[j]$ output by $\mathcal{A}$ are within $2\alpha$ of the inner products $\langle q^{(j)},y\rangle$, and so $\mathcal{A}$ is $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$. ∎

### v3

@@ 3.2 Adaptively estimating a good flippancy bound $w$
In this section, we leverage the privacy and accuracy guarantees of Algorithm 1 to construct a new mechanism (Algorithm 3) for estimating $\mathsf{CountDistinct}$. It achieves the privacy and accuracy guarantees of Theorem 3.1, when the maximum flippancy is not known upfront. Algorithm 3 instantiates $\log T+1$ different copies $\mathcal{B}_{0},\dots\mathcal{B}_{\log T}$ of Algorithm 1 with flippancy bounds $2^{0},\dots,2^{\log T}$, respectively (the maximum flippancy of a stream is at most $T$). To obtain an accurate estimate of the distinct elements count, at each time $t\in[T]$, we privately select $i\in[0,\log T]$ such that the output of $\mathcal{B}_{i}$ satisfies the desired accuracy guarantee for the stream entries $x[1:t]$ received so far. Selecting such $i$ amounts to selecting a good bound on the maximum flippancy of the stream $x[1:t]$. Next, we describe how to obtain this bound using the sparse vector technique (Algorithm 2).
The maximum flippancy has high sensitivity: changing one stream entry can drastically change the maximum flippancy. However, the number of items with flippancy greater than any particular threshold is a function of sensitivity one. Furthermore, since Algorithm 1 when run with flippancy bound $w$ already has error about $\sqrt{w/\rho}$, its accuracy guarantee remains asymptotically the same even if it simply ignores that many elements with flippancy greater than $w$. Thus, Algorithm 3 uses the sparse vector technique to maintain an upper bound on the flippancy of $x[1:t]$ such that not too many elements in $x[1:t]$ violate that bound. This bound, in combination with the error guarantee of Algorithm 1, suffices to provide the desired low error guarantee. Since the sparse vector algorithm remains differentially private even when its queries are chosen adaptively, the privacy guarantees of Algorithm 3 follow from the privacy of Algorithms 1 and 2.
@@
8:         Return $\mathsf{Above}$
@@
11:         Return $\mathsf{Below}$
@@
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, stream $x\in\mathcal{U}_{\pm}^{T}$
@@
10:     for all $i\in[0,\log{T}]$ do
11:         Send $x[t]$ to mechanism $\mathcal{B}_{i}$ and get output $s_{i,t}$
@@
16:     Return $s_{j,t}$ for $j=\log(w_{\max}[t])$triangleright Note that $j\in[0,\log T]$
@@
Algorithm 2 is $\sqrt{2\rho}$-DP (and therefore $\rho$-zCDP). Let $k$ be the index of the last “$\mathsf{Above}$” query answered by Algorithm 2 (before cutoff $c$ has been crossed). For all $\beta\in(0,1)$, with probability at least $1-\beta$, the vector of answers to the queries $q_{1},\dots,q_{k}$ is $\gamma$-accurate for $\gamma=\frac{8c(\ln k+\ln(2c/\beta))}{\sqrt{2\rho}}$.
@@
We start by showing that Algorithm 3 is $\rho$-item-level-zCDP. Algorithm 3 accesses the stream $x$ via Algorithm 2 and the algorithms $\mathcal{B}_{i}$ for $i\in[0,\log T]$ (instantiations of Algorithm 1). By Theorem 3.2, Algorithm 1 with privacy parameter $\rho^{\prime}$ is $\rho^{\prime}$-item-level-zCDP. Since we use $(\log T+1)$ instantiations of Algorithm 1, each with privacy parameter $\frac{\rho}{2(\log T+1)}$, by composition, the aggregate of the calls to Algorithm 1 is $(\frac{\rho}{2})$-item-level-zCDP. We now show that the aggregate of the calls to Algorithm 2 is $(\frac{\rho}{2})$-item-level-zCDP. Note that the queries $q_{t}$ for $t\in[T]$ considered in Step 13 of Algorithm 3 have sensitivity $1$ for item-neighboring streams (the number of items with flippancy above a certain threshold can change by at most 1 for item-neighboring streams). By Theorem 3.7, the aggregate of the calls to Algorithm 2 is $(\frac{\rho}{2})$-item-level-zCDP. Another invocation of the composition lemma gives that Algorithm 3 is $\rho$-item-level-zCDP.
We now analyze the accuracy of Algorithm 3. Set $\beta_{\mathsf{SVT}}=0.005$, $k=T$, $c=\log T$, and $\gamma_{\mathsf{SVT}}=\frac{8\log T(\log T+\log(400\log T))}{\sqrt{2\rho}}$. Let $E$ be the event that the vector of answers output by the sparse vector algorithm (Algorithm 2) until the cutoff point $\log T$ is $\gamma_{\mathsf{SVT}}$-accurate. By Theorem 3.7, $\Pr[E]\geq 0.996$. We condition on $E$ for most of the following proof.
@@
Let $w_{t}$ be the (true) maximum flippancy of the sub-stream $x[1:t]$, consisting of the first $t$ entries of the input stream $x\in\mathcal{U}_{\pm}^{T}$ to Algorithm 3. Then, for all $t\in[T]$, when the algorithm reaches Step 16,
@@
Let $z$ be the value of $w_{\max}[t]$ when Algorithm 3 reaches Step 16. If $z=1$ then $z=w_{\max}[t]\leq 2\gamma^{2}_{\mathsf{SVT}}$ since $T>1,\rho<1$. So, instead assume that $z\geq 2$. Let $t^{*}\leq t$ be the time step where $w_{\max}[t^{*}]$ is doubled from $\frac{z}{2}$ to $z$ during an execution of Step 15 of the $\mathbf{while}$ loop. This only happens if Algorithm 2 outputs $\mathsf{Above}$ for the following query:
@@
Suppose there is some $t\in[T]$ during which $\mathsf{count}\geq c$. Consider the last time step $t^{*}\in[T]$ when Step 9 of Algorithm 2 is run (for this time step, $\mathsf{count}=c-1$). At this time step, $w_{max}[t^{*}]$ doubles from $\frac{T}{2}$ to $T$, after which it never changes again. By case (1), we have that $w_{max}[t^{*}]=T\leq\max(2w_{t^{*}},2\gamma^{2}_{\mathsf{SVT}})$. Since for all $t\geq t^{*}$ it holds $w_{t^{*}}\leq w_{t}$ and $w_{max}[t]=w_{max}[t^{*}]$, then $w_{max}[t]\leq\max(2w_{t},2\rho\gamma^{2}_{\mathsf{SVT}})$ for all $t\geq t^{*}$. This concludes the proof of Claim 3.9. ∎
Now we substitute the upper bound on $w_{max}[t]$ from Claim 3.9 into Equation (4) and apply $w_{t}\leq w$. We get that, for sufficiently large $T$, conditioned on event $E$, with probability at least $0.995$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is
@@
Finally, by a union bound over the event $E$ and the event that the error of Algorithm 3 is greater than Equation (5) we obtain: For sufficiently large $T$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is bounded by Equation (5) with probability at least $0.99$. ∎
@@
Let $f\colon\mathcal{U}_{\pm}^{*}\to\mathbb{R}$ be a function of $\ell_{2}$-sensitivity at most $1$. Define $F\colon\mathcal{U}_{\pm}^{T}\to\mathbb{R}^{T}$ so that $F(x)=[f(x[1:1]),\dots,f(x[1:T])]$. Fix $\rho\in(0,1]$ and sufficiently large $T\in\mathbb{N}$. Then, there exists a $\rho$-item-level-zCDP mechanism for estimating $F$ in the continual release model that is $\alpha$-accurate where $\alpha=\operatorname{O}\left(\min\left\{\sqrt[3]{\frac{T\log T}{\rho}},T\right\}\right).$
Note that $\mathsf{CountDistinct}(x)[t]$ has $\ell_{2}$-sensitivity one for item-neighboring streams for all $t\in[T]$. Let $\mathcal{M}^{\prime}$ be the mechanism from Theorem 3.10. Then $\mathcal{M}^{\prime}$ can be used for estimating $\mathsf{CountDistinct}$ under continual release for turnstile streams with the error guarantee stated in Theorem 3.10. When the maximum flippancy of the stream is larger than roughly $\rho^{1/3}T^{2/3}$, the mechanism $\mathcal{M}^{\prime}$ achieves better error than that of Theorem 3.1 (and it achieves worse error when the maximum flippancy of the stream is smaller than this threshold). A simple modification of Algorithm 3 can get the best of both worlds—instead of having base mechanisms $\mathcal{B}_{0},\dots,\mathcal{B}_{\log T}$ that each run Algorithm 1 with different flippancy parameters as input, we only have $k+2=\mathsf{min}(\rho^{1/3}T^{2/3},T)$ base mechanisms $\mathcal{B}_{0},\dots,\mathcal{B}_{k+1}$. Out of these, $\mathcal{B}_{0},\dots,\mathcal{B}_{k}$ run Algorithm 1, whereas $\mathcal{B}_{k+1}$ runs $\mathcal{M}^{\prime}$. This modified algorithm has error $O\left(\mathsf{min}\left(\sqrt{\frac{w}{\rho}}\operatorname{polylog}T,\sqrt[3]{\frac{T\log T}{\rho}},T\right)\right)$. The proof is similar to the proof of Theorem 3.1, with the only difference that for analyzing the error of base mechanism $\mathcal{B}_{k+1}$ we use the error guarantee of the recompute-mechanism $\mathcal{M}^{\prime}$.
@@ 4 Event-level privacy lower bound
In this section, we prove Theorem 1.6, providing a strong lower bound on the parameter $\alpha$ for every $\alpha$-accurate, event-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
@@ 4.1 Reduction from $\mathsf{InnerProducts}$
We obtain our lower bound by showing that every mechanism for $\mathsf{CountDistinct}$ for turnstile streams can be used to obtain an algorithm with similar accuracy guarantees for $\mathsf{InnerProducts}$, the problem of estimating answers to inner product queries in the batch model. The reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$ combines two ideas: one is the sequential embedding technique introduced by Jain et al. [44] to prove lower bounds in the continual release model and the other is a connection between the inner product of two vectors and the number of distinct elements in the concatenation of two corresponding streams. The latter idea was used by Mir et al. [50] to give lower bounds for pan-private algorithms for counting the number of distinct elements. The reduction is presented in Algorithm 4. With this reduction, we then use previously established lower bounds on accuracy for $\mathsf{InnerProducts}$ [21, 23, 50, 17] to obtain our lower bound on $\mathsf{CountDistinct}$. We start by proving Lemma 4.4 (the reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$). In Section 4.2, we use Lemma 4.4 to complete the proof of Theorem 1.6.
Algorithm 4 crucially uses the following connection between the inner product of two vectors and the number of distinct elements in the concatenation of two corresponding streams.
Definition 4.1 (Stream indicator).
Remark 4.2.
For every pair of insertion-only streams $x$ and $x^{\prime}$,
@@
2:Output: Estimates of inner products $b=(b[1],\dots,b[k])\in\mathbb{R}^{k}$
3:Define the universe $\mathcal{U}=[n]$
4:Initialize stream $z^{(0)}=\bot^{n}$
@@
6:     If $y[i]=1$ set $z^{(0)}[i]=+i$
7:Initialize streams $z^{(1)}=\perp^{2n},\dots,z^{(k)}=\perp^{2n}$ and a vector $r$ of length $(2k+1)n$
@@
9:     Set $z^{(j)}[i]=+i$ and $z^{(j)}[n+i]=-i$
10:Run $\mathcal{M}$ on the stream $x\leftarrow z^{(0)}\circ z^{(1)}\circ z^{(2)}\circ\dots\circ z^{(k)}$ and record the answers as vector $r$
11:for all $j\in[k]$ do
12:     Compute $\|q^{(j)}\|_{0}$ and let $b[j]=\|q^{(j)}\|_{0}+r[n]-r[2jn]$
13:Return the estimates $(b[1],\dots,b[k])$
@@
We now show that if the input mechanism $\mathcal{M}$ to Algorithm 4 is accurate for $\mathsf{CountDistinct}$, then Algorithm 4 is accurate for $\mathsf{InnerProducts}$.
@@
Let $\mathcal{A}$ be Algorithm 4. For all $\varepsilon>0,\delta\in[0,1),\alpha\in\mathbb{R}^{+}$ and $n,T,k\in\mathbb{N}$, where $T\geq(2k+1)n$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-event-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $2k$, then batch algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$.
@@
Algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-event-level-DP because $\mathcal{M}$ is $(\varepsilon,\delta)$-event-level-DP and changing a record of the dataset $y$ corresponds to changing a single entry of the stream $x$, and more specifically, an entry of the stream $z^{(0)}$ constructed in Step 6 of Algorithm 4.
We are left to prove the accuracy of $\mathcal{A}$. Fix queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$. First, observe that $z^{(0)}$ is constructed so that $y$ is its stream indicator vector. Similarly, observe that for all $j\in[k]$ , the stream $z^{(j)}$ is constructed so that $q^{(j)}$ is the indicator vector for $z^{(j)}[1:n]$, namely, the first half of $z^{(j)}$.
Next, since at time $2jn$, all of the stream entries pertaining to earlier queries $q^{(1)},\dots,q^{(j-1)}$ have been deleted and those pertaining to $q^{(j)}$ have been inserted, $\|h_{x[1:2jn]}\|_{0}=\|h_{z^{(0)}\circ z^{(j)}[1:n]}\|_{0}$ for $j\in[k]$.
The streams $z^{(0)}$ and $z^{(j)}[1:n]$ for $j\in[k]$ are all insertion-only streams. By Remark 4.2,
$\langle h_{z^{(0)}},h_{z^{(j)}[1:n]}\rangle=\|h_{z^{(0)}}\|_{0}+\|h_{z^{(j)}[1:n]}\|_{0}-\|h_{z^{(0)}\circ z^{(j)}[1:n]}\|_{0}.$
As observed earlier, $h_{z^{(0)}}=y$, $h_{z^{(j)}[1:n]}=q^{(j)}$, and $\|h_{x[1:2jn]}\|_{0}=\|h_{z^{(0)}\circ z^{(j)}[1:n]}\|_{0}$. Thus,
$\langle y,q^{(j)}\rangle$ $=\|h_{z^{(0)}}\|_{0}+\|q^{(j)}\|_{0}-\|h_{x[1:2jn]}\|_{0}.$ (6)
@@
Since the mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ is $\alpha$-accurate on the constructed stream then, with probability at least $0.99$, the answers of $\mathcal{M}$ are within additive error $\alpha$ of the distinct counts of the corresponding stream prefixes. Condition on this event for the rest of this proof. Then, $|r[n]-\|h_{z^{(0)}}\|_{0}|\leq\alpha$. Similarly, $|r[2jn]-\|h_{x[1:2jn]}\|_{0}|\leq\alpha$ for all $j\in[k]$. Additionally, $\|q^{(j)}\|_{0}$ is computed exactly by $\mathcal{A}$. Hence, by the triangle inequality, Equation 6, and the setting of $b[j]$ in Step 12, we have that $|b[j]-\langle y,q^{(j)}\rangle|\leq 2\alpha$ for all $j\in[k]$. Hence, with probability at least $0.99$, all of the estimates $b[j]$ returned by $\mathcal{A}$ are within $2\alpha$ of the inner products $\langle q^{(j)},y\rangle$, and so $\mathcal{A}$ is $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$. ∎